OpenStudy (anonymous):

Write the inverse variation equation of the following: Y varies inversely with x and y = -12 when x = 6. xy = -2 xy = -72 xy = -0.5

OpenStudy (anonymous):

y = k/x and then plug in the given values.

OpenStudy (anonymous):

-72?

OpenStudy (anonymous):

-12 = k/6 Now multiply both sides by 6 6(-12) = k So yes... k = -72 y = -72/x or xy = -72